OpenStudy (anonymous):

Compute the monthly payments for an add-on interest loan of $2,000, with an annual interest rate of 18 percent and a term of 3 years. Round to the nearest cent as needed.

OpenStudy (anonymous):

well first of all thats an insane interest rate haha but anyway, here's how you do it...

OpenStudy (anonymous):

hahaha i mean its what my math teacher decided to throw out there

OpenStudy (anonymous):

is this stuff like easy for you because i struggle? ha

OpenStudy (anonymous):

FV=($2000)(1+.18)^3=$3286.06 So you owe a total of $3286.06 by the end of the 3 years. Now splitting that up in to monthly payments over the 3 years: $3286.06/(12 months * 3 years) = 3286.06/36 = $91.28 So your monthly payments would be $91.28.

OpenStudy (anonymous):

well i guess after awhile of doing it, it gets easier. but at first it can definitely be confusing

OpenStudy (anonymous):

i cant believe your just on here helping! haha

OpenStudy (anonymous):

yeah i guess im in a giving mood today

OpenStudy (anonymous):

...or maybe im just procrastinating really hard on my finals haha

OpenStudy (anonymous):

i mean im not mad at it haha

OpenStudy (anonymous):

no you've got a good understanding it seems like